OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (anonymous):

1) 4x - y = 1 2) 4x + y = 1 3) x - 4y = -13 4) 4x - y = -7 Save